Circular No. 6200

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) M. Hicks, Lunar and Planetary Laboratory, reports the following JHK

photometry, obtained on July 27 with the 1.8-m telescope at the Lowell Observatory’s Anderson Mesa Station, using the OSIRIS infrared system in imaging mode. Synthetic apertures of diameter 12′′, 18′′ and 24′′ were applied to obtain the following magnitudes: J(12) = 12.22 ± 0.07, J(18) = 12.02 ± 0.07, J(24) = 11.87 ± 0.09; H(12) = 11.93 ± 0.02, H(18) = 11.66 ± 0.03, H(24) = 11.51 ± 0.12; K(12) = 11.86 ± 0.05, K(18) = 11.56 ± 0.05, K(24) = 11.37 ± 0.07. No color differences with respect to aperture size were found, and the J − H and J − K magnitudes are consistent with the solar spectrum, suggesting optically dominant grain size ≥ 1–2 μm, typical for comets.

RX J1802.1+1804 J. Greiner, Max-Planck-Institute for Extraterrestrial Physics; R. Re-

millard, Center for Space Research, Massachusetts Institute of Technology; and C. Motch, Strasbourg Observatory, report in response to IAUC 6195: “The new AM Her system RX J1802.1+1804 was discovered in ROSAT data as a highly variable and soft x-ray source. Follow-up pointed observations revealed a periodic flux modulation with the ephemeris T (HJD) = 2448907.26986597+0.078498(115)E and including a sharp and nearly complete eclipse of the x-ray emission. The x-ray spectrum averaged over the full period is dominated by soft emission below 0.5 keV. A blackbody plus a thermal bremsstrahlung model with a fixed temperature of 10 keV results in kT_{bb} = 30 ± 20 eV. The mean unabsorbed blackbody flux (excluding the eclipse) is 8 × 10^{−11} erg cm^{−2} s^{−1} in the ROSAT band—the number given on IAUC 6195 is wrong—and and the ratio of soft (blackbody) to hard (bremsstrahlung) photon flux (integrated over 0.1–2.4 keV) is 100. RX J1802.1+1804 is optically identified with an object of V = 15. The optical spectrum is typical for a high-excitation cataclysmic variable with strong emission lines of H, He I and He II. Phase-resolved spectroscopy shows a strong velocity wave, giving a best-fit period of 1.8841 ± 0.0017 hr. Photometric observations have revealed 0.1-mag variations, but no eclipse. The evidence of an x-ray eclipse without a corresponding optical eclipse implies that the x-ray-emitting region is confined to a small region on the whitedwarf surface. These and more details have been presented at the Abano-Terme Conference on Cataclysmic Variables and Related Objects in June 1994; preprints are available on request from jcg@mpe-garching.mpg.de.”

1995 August 3 Brian G. Marsden